# Robust aluminum nitride passivation of silicon carbide with near-surface spin defects

### v1

@@ Introduction
The singly charged defect at the interior site, S1, has a total of five electrons in the dangling bonds left behind by the monovacancy. It is a spin-3/2 defect, just as in bulk SiC Joshi and Dev (2022); Carter et al. (2015); Soykal et al. (2016); Economou and Dev (2016). Figures 4(a)-(d) give the details of the electronic structure for this defect. Figure 4(a) is a plot of TDOS (in gray) of the defective SiC/AlN-H NW, along with the projected DOS contributed by the SiC core (in gold), AlN shell (in thin black line) and the four NN carbon-atoms, ’s surrounding the defect (in blue). The plot clearly shows that the defect creates localized defect states in the band gap. The DOS contributed by SiC shows that the defect states are contributed by the core with negligible contribution from the AlN shell. The DOS contribution from the four ’s and the isosurface plot for the spin density [] in Fig. 4(b) further shows that the most of the spin-3/2 of the defect comes from the unpaired electrons in the dangling bonds associated with the carbon atoms surrounding the defect. To understand the origin of spin-3/2 for the defect, we show the energy-level diagram in Fig. 4(c). The defective NW structure has the symmetry, unlike the symmetry for the defect in the bulk Joshi and Dev (2022); Sayou Ngomsi et al. (2024), and thus, all of the defect states are singlets. The five electrons in the dangling bonds fill the “atomic-like” localized defect states such that there are three more electrons in the spin-up channel (majority-spin) than the spin-down channel (minority spin), resulting in the spin-3/2 for the defect.
@@
The optical excitation happens between the filled and the empty defect states in the spin-down channel. Figure 4(d) shows the charge density plots for these optically-active empty and filled states. The calculated value of the ZPL is 1.12 eV, which differs from the bulk ZPL-value of 1.19 eV by about 75 meV. As discussed in earlier works Joshi and Dev (2022); Sayou Ngomsi et al. (2024), the difference in ZPL for a defect close to the surfaces and/or interfaces can be attributed to several inter-connected factors, with the most significant being: (i) local strains, (ii) differences in Stokes shift, which result from different structural relaxation around the defect upon photoexcitation, and (iii) different placements and hence gaps between the optically-active defect states, resulting in different absorption energies. To estimate the strain-induced change in the ZPL, we used the experimentally-determined strain-coupling parameter, 1130 meV(strain, for strain in the basal direction Vásquez et al. (2020). The bond lengths along the basal plane at the S1-site are 0.39% longer as compared to the ideal basal distance. This yields a ZPL-change of 4.38 meV, showing that strain is relatively less important in changing the ZPL. The Stokes shift, given by in Fig. 1(c), is calculated to be 122.3 meV for the defect at S1, while it is 104.9 meV for the defect in the bulk material. The 17.4 meV difference in the Stokes shift accounts for some of the calculated ZPL shift. In the case of at the S1-site in the SiC/AlN-H NW, the largest contributor to the ZPL-shift is from the differences in vertical absorption energies [ in Fig. 1(c)] for the defect at the S1 site (1.24 eV) and for the bulk (1.30 eV), with a 57.89 meV decrease in the calculated value of for the defect in core-shell NW as compared to the bulk. To see the influences of the AlN-shell on the optical properties of the interior defect, we also calculated the optical properties of at an interior-site in the SiC-H NW. We find that the ZPL for this defect (1.16 eV) is much closer to that for the bulk SiC, with a 30 meV difference in the ZPL. For the monovacancy in the interior site of SiC-H NW, the contributions to the ZPL-shift from strain, changes in Stokes shift and the vertical excitation are found to be 1.57 meV, 11.77 meV and 16.87 eV, respectively. Hence, the largest change produced by AlN shell is in the placement of the defect states within the band gap (as compared to a SiC-H NW), owing to the changes in the electronic structure properties upon creating a core-shell NW.
.2.2 at the S2 (interface) site in SiC/AlN-H NW
@@
We first concentrate on explaining the calculated electronic structure to explain the spin-1/2 of the defect and then tackle the question: “ where is the sixth electron?” In Fig. 5(a), we plot the total and the partial DOS plots for the spin-1/2 defect at the interface, showing that the defect introduces very localized (sharp) defect states in the band gap. The DOS contributions from the and orbitals of the NN carbons atoms (in blue) and nitrogen atom (in green) show that the in-gap sharp states originate in the dangling bonds on the surrounding atoms. The energy-level diagram is shown in Fig. 5(b). The defect states resonant with VB are shown as “fuzzy” lines as these defect states hybridize with the states in the VB and their position can no longer be uniquely given. A comparison with the energy level diagram for the interior defect [Fig. 4(c)] shows that the three nearly-degenerate defect states of the interior defect have undergone larger crystal field splitting at the interface, with the distribution of the five electrons resulting in net spin-1/2 for the interfacial defect. The spin density for the defect at the interface is shown in Figure 5(c). Pink (green) color corresponds to positive (negative) isovalues. The magnetic moment of the axial NN carbon is antiferromagnetic aligned to the moments of the NN carbons along the basal plane, resulting in the net spin-1/2 for the defect. The close-up of the spin density plot (side-view) shows only the three atoms and atom, along with the Al atoms bonded to the -atom. The three Al atoms are shown to highlight that, although nominally a negatively-charged NV center, this defect has a different chemical environment as compared to the same defect in SiC bulk or SiC NW without an AlN shell. To demonstrate that the antiferromagnetic alignment is not a result of proximity to the surface, we created a negatively charged defect center (subsequently referred to as the “NV-center”) in a SiC-H NW at the same site. We find that the near-surface, negatively-charged NV center in SiC-H NW is a spin-1 defect (similar to the NV-center in the bulk SiC) with a spin-density distribution shown in Fig. 5(d). Hence, neither the nanostructuring of SiC, nor the proximity to the surface are responsible for the behavior of the interfacial defect in SiC/AlN-H NW.
@@
In order to answer the question posed in the previous paragraph (“where is the sixth electron?”), we calculated the charge-density differences between the singly-charged and neutral defects [] for three systems: at the S2 (interface) site in SiC/AlN-H NW (effectively, an NV center), at the S1 (interior) site in SiC/AlN-H NW, and an NV center in a SiC-H NW. The charge densities for the neutral charge states of all three defects were calculated with the respective equilibrium geometries obtained for the three negatively-charged structures. Figures 6(a)-(c) are the plots of the charge density differences for the three systems. Yellow (blue) color corresponds to regions with charge accumulation (depletion). Overall, Figs 6(a)-(c) show the sense in which charge is redistributed upon addition of an electron to the respective neutral defects. The charge density differences for the monovacancy at S2 in Fig. 6(a) shows that the electron added to the neutral defect at S2 does not get localized at the defect site. It is instead distributed over a large region, including the AlN shell. One may suspect that this is either due to the nanostructuring of SiC itself or due to the proximity of the defect to the surface. That this is not the case can be seen in the highly-localized charge redistribution upon the addition of an electron for the other two structures shown in Figs 6(b) and (c). Hence, we find that the defect at S2 is fundamentally different from the other two defects and we attribute this to the different chemical environment of the defect at the AlN-SiC interface. Interestingly, the interfacial defect acquires a spin-1 ground state when it is doubly-negatively charged. This charge state becomes stable for the at the S2 site for highly negatively doped NW as seen in Fig. 3(b). In our nanostructure, both and charge states of the interfacial defect have a spin-active ground state, with the magnetic structures being more stable than the non-magnetic structures by 387.03 meV and 415.29 meV, respectively. The ZPLs for the singly and doubly charged monovacancies at the S2-site are 0.68 eV and 0.77 eV, respectively. The latter is comparable to the ZPL of 0.84 eV for the spin-1 -center (with N substituting a basal carbon) in bulk 2H-SIC. The difference in the ZPLs of the two spin-1 defects originates mostly from the much larger Stokes shift for the doubly-charged interfacial monovacancy (158.48 meV) as compared to that for the singly-charged -center in bulk 2H-SIC (82.23 meV).
Conclusion
@@
The loss of photostability of near-surface, defect-based quantum emitters within nanostructured semiconductors is a severe bottleneck for their successful deployment in different quantum technologies. In this work, we have demonstrated that the core-shell structure, which consists of a SiC core passivated with an AlN shell, can effectively remove the detrimental surface states from the band gap of the nanostructured SiC host, elucidating the findings of an earlier experimental work Polking et al. (2018). In such a core-shell model, the AlN shell replaces simple atomic/molecular adsorbates, which will otherwise be used to passivate the surface states but are limited by their lack of long-term chemical and thermal stability. The presence of an AlN shell not only resolves the issue of charge-state instability of the near-surface defect, it also restores the optical properties of this spin-3/2 defect. In addition, we identify a -related defect at the interface of the SiC core and the AlN shell. This interfacial defect is nominally similar to the NV-center in SiC (with nitrogen along the basal plane), but displays different spin and optical properties from an NV-center in SiC due to its distinct chemical environment at the interface. Our atomistic study of defects in SiC/AlN-H NW shows a promising means of resolving the optical stability issues of quantum emitters in nanostructured host semiconductors via creation of core-shell structures or growth/deposition of an appropriate semiconductor as a passivation layer on thin films of host semiconductors.
Acknowledgements.
@@
Acknowledgements: We acknowledge support by the National Science Foundation under NSF grant number DMR-1738076 and the STC Center for Integrated Quantum Materials under NSF Grant number DMR-1231319. This work used the Expanse and Bridges2 clusters at SDSC and PSC, respectively, through allocation PHY180014 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants No. 2138259, No. 2138286, No. 2138307, No. 2137603, and No. 2138296.Author Contributions: P.D. conceived and directed the study. C.A.S.N. and P.D. performed the calculations. P.D. wrote the final manuscript with inputs from the first draft written by C.A.S.N. Both authors proof-read and reviewed the final revision of the manuscript.

### v2

@@ 1.2 Silicon vacancy in passivated SiC/AlN-H NW
The singly charged defect at the interior site, S1, has a total of five electrons in the dangling bonds left behind by the monovacancy. It is a spin-3/2 defect, just as in bulk SiC 22, 7, 50, 51. Figures 4(a)-(d) give the details of the electronic structure for this defect. Figure 4(a) is a plot of TDOS (in gray) of the defective SiC/AlN-H NW, along with the projected DOS contributed by the SiC core (in gold), AlN shell (in thin black line) and the four NN carbon-atoms, ’s surrounding the defect (in blue). The plot clearly shows that the defect creates localized defect states in the band gap. The DOS contributed by SiC shows that these defect states are contributed by the core with negligible contribution from the AlN shell. The DOS contribution from the four ’s and the isosurface plot for the spin density [] in Figure 4(b) further shows that the most of the spin-3/2 of the defect comes from the unpaired electrons in the dangling bonds associated with the carbon atoms surrounding the defect. The localized nature of the defect states itself can be traced back to the highly localized nature of carbon’s -hybridized and orbitals that form the dangling bonds. The large exchange interactions between the unpaired electrons in the nearly-dispersionless defect states results in the spin-splitting of the defect states seen in Figure 4(a) 52, 53, 54.
To understand the origin of spin-3/2 for the defect, we show the energy-level diagram in Figure 4(c). The defective NW structure has the symmetry, unlike the symmetry for the defect in the bulk 22, 23, and thus, all of the defect states are singlets. The five electrons in the dangling bonds fill the “atomic-like" localized defect states such that there are three more electrons in the spin-up channel (majority-spin) than the spin-down channel (minority spin), resulting in the spin-3/2 for the defect. The PH-passivated SiC/AlN NW gives identical results for the defect at the S1-site [see Supplementary Figure S5].
@@
The optical excitation happens between the filled and the empty defect states in the spin-down channel. Figure 4(d) shows the charge density plots for these optically-active empty and filled states. The calculated value of the ZPL is 1.12 eV, which differs from the bulk ZPL-value of 1.19 eV by about 75 meV. As discussed in earlier works 22, 23, the difference in ZPL for a defect close to the surfaces and/or interfaces can be attributed to several inter-connected factors, with the most significant being: (i) local strains, (ii) differences in Stokes shift, which result from different structural relaxation around the defect upon photoexcitation, and (iii) different placements and hence gaps between the optically-active defect states, resulting in different absorption energies. To estimate the strain-induced change in the ZPL, we used the experimentally-determined strain-coupling parameter, 1130 meV(strain, for strain in the basal direction 17. The bond lengths along the basal plane at the S1-site are 0.39% longer as compared to the ideal basal distance. This yields a ZPL-change of 4.38 meV, showing that strain is relatively less important in changing the ZPL. The Stokes shift, given by in Figure 1(c), is calculated to be 122.3 meV for the defect at S1, while it is 104.9 meV for the defect in the bulk material. The 17.4 meV difference in the Stokes shift accounts for some of the calculated ZPL shift. In the case of at the S1-site in the SiC/AlN-H NW, the largest contributor to the ZPL-shift is from the differences in vertical absorption energies [ in Figure 1(c)] for the defect at the S1 site (1.24 eV) and for the bulk (1.30 eV), with a 57.89 meV decrease in the calculated value of for the defect in core-shell NW as compared to the bulk. To see the influences of the AlN-shell on the optical properties of the interior defect, we also calculated the optical properties of at an interior-site in the SiC-H NW. We find that the ZPL for this defect (1.16 eV) is much closer to that for the bulk SiC, with a 30 meV difference in the ZPL. For the monovacancy in the interior site of SiC-H NW, the contributions to the ZPL-shift from strain, changes in Stokes shift and the vertical excitation are found to be 1.57 meV, 11.77 meV and 16.87 meV, respectively. Hence, the largest change produced by AlN shell is in the placement of the defect states within the band gap (as compared to a SiC-H NW), owing to the changes in the electronic structure properties upon creating a core-shell NW.
1.2.2 at the S2 (interface) site in SiC/AlN-H NW
@@
We first concentrate on explaining the calculated electronic structure to explain the spin-1/2 of the defect and then tackle the question: “ where is the sixth electron?” In Figure 5(a), we plot the total and the partial DOS plots for the spin-1/2 defect at the interface, showing that the defect introduces very localized (sharp) defect states in the band gap. The DOS contributions from the and orbitals of the NN carbons atoms (in blue) and nitrogen atom (in green) show that the in-gap sharp states originate in the dangling bonds on the surrounding atoms. The energy-level diagram is shown in Figure 5(b). The defect states resonant with VB are shown as “fuzzy” lines as these defect states hybridize with the states in the VB and their position can no longer be uniquely given. A comparison with the energy level diagram for the interior defect [Figure 4(c)] shows that the three nearly-degenerate defect states of the interior defect have undergone larger crystal field splitting at the interface, with the distribution of the five electrons resulting in net spin-1/2 for the interfacial defect. We obtain a qualitatively similar energy level diagram for the S2-site defect in the SiC/AlN-PH NW [see Supplementary Figure S6 and Supplementary Note-II].
The spin density for the defect at the interface is shown in Figure 5(c). Pink (green) color corresponds to positive (negative) isovalues. The magnetic moment of the axial NN carbon is antiferromagnetic aligned to the moments of the NN carbons along the basal plane, resulting in the net spin-1/2 for the defect. The close-up of the spin density plot (side-view) shows only the three atoms and atom, along with the Al atoms bonded to the -atom. The three Al atoms are shown to highlight that, although nominally a negatively-charged NV center, this defect has a different chemical environment as compared to the same defect in SiC bulk or SiC NW without an AlN shell. To demonstrate that the antiferromagnetic alignment is not a result of proximity to the surface, we created a negatively charged defect center (subsequently referred to as the “NV-center”) in a SiC-H NW at the same site. We find that the near-surface, negatively-charged NV center in SiC-H NW is a spin-1 defect (similar to the NV-center in the bulk SiC) with a spin-density distribution shown in Figure 5(d). Hence, neither the nanostructuring of SiC, nor the proximity to the surface are responsible for the behavior of the interfacial defect in SiC/AlN-H NW.
@@
In order to answer the question posed in the previous paragraph (“where is the sixth electron?"), we calculated the charge-density differences between the singly-charged and neutral defects [] for three systems: at the S2 (interface) site in SiC/AlN-H NW (effectively, an NV center), at the S1 (interior) site in SiC/AlN-H NW, and an NV center in a SiC-H NW. The charge densities for the neutral charge states of all three defects were calculated with the respective equilibrium geometries obtained for the three negatively-charged structures. Figures 6(a)-(c) are the plots of the charge density differences for the three systems. Yellow (blue) color corresponds to regions with charge accumulation (depletion). Overall, Figs 6(a)-(c) show the sense in which charge is redistributed upon addition of an electron to the respective neutral defects. The charge density differences for the monovacancy at S2 in Figure 6(a) shows that the electron added to the neutral defect at S2 does not get localized at the defect site. It is instead distributed over a large region, including the AlN shell. One may suspect that this is either due to the nanostructuring of SiC itself or due to the proximity of the defect to the surface. That this is not the case can be seen in the highly-localized charge redistribution upon the addition of an electron for the other two structures shown in Figs 6(b) and (c). Hence, we find that the defect at S2 is fundamentally different from the other two defects and we attribute this to the different chemical environment of the defect at the AlN-SiC interface. Interestingly, the interfacial defect acquires a spin-1 ground state when it is doubly-negatively charged. This charge state becomes stable for the at the S2 site for highly negatively doped NW as seen in Figure 3(b). In our nanostructure, both and charge states of the interfacial defect have a spin-active ground state, with the magnetic structures being more stable than the non-magnetic structures by 387.03 meV and 415.29 meV, respectively. The ZPLs for the singly and doubly charged monovacancies at the S2-site are 0.68 eV and 0.77 eV, respectively. The latter is comparable to the ZPL of 0.84 eV for the spin-1 -center (with N substituting a basal carbon) in bulk 2H-SiC. The difference in the ZPLs of the two spin-1 defects originates mostly from the much larger Stokes shift for the doubly-charged interfacial monovacancy (158.48 meV) as compared to that for the singly-charged -center in bulk 2H-SIC (82.23 meV).
Conclusions
@@
The loss of photostability of near-surface, defect-based quantum emitters within nanostructured semiconductors is a severe bottleneck for their successful deployment in different quantum technologies. In this work, we have demonstrated that the core-shell structure, which consists of a SiC core passivated with an AlN shell, can effectively remove the detrimental surface states from the band gap of the nanostructured SiC host, elucidating the findings of an earlier experimental work 27. In such a core-shell model, the AlN shell replaces simple atomic/molecular adsorbates, which will otherwise be used to passivate the surface states but are limited by their lack of long-term chemical and thermal stability. The presence of an AlN shell not only resolves the issue of charge-state instability of the near-surface defect, it also restores the optical properties of this spin-3/2 defect. In addition, we identify a -related defect at the interface of the SiC core and the AlN shell. The interfacial defect is nominally similar to the NV-center in SiC (with nitrogen along the basal plane), but displays different spin and optical properties from an NV-center in SiC due to its distinct chemical environment at the interface. Since this defect exists at the interface where other factors, such as strains and other defect complexes may be present, further careful investigations are needed to determine its usefulness under different conditions. Our atomistic study of defects in the SiC/AlN-H NW shows a promising means of resolving the optical stability issues of quantum emitters in nanostructured host semiconductors via creation of core-shell structures or growth/deposition of an appropriate semiconductor as a passivation layer on thin films of host semiconductors. Lastly, as noted in our previous work 23, although we have employed a small-diameter nanowire in our calculations, the primary findings of this study are robust and should be applicable to the near-surface defects in larger nanostructures.
2 Associated Content
Data Availability Statement
The crystal structures data that support the findings of this study are available
from the corresponding author upon reasonable request.
Supporting Information
The Supporting Information is available free of charge.
Supplementary figures and notes contrasting the resulting density of states, defect formation energies and the electronic structure of the defects in core shell nanowires, which is passivated with either real or pseudohydrogens.
3 Author Information
Corresponding Author
Pratibha Dev – Department of Physics and Astronomy, Howard University, Washington, D.C. 20059, USA, Laboratory for Physical Sciences, College Park, Maryland 20740, United States; orcid.org/0000-0002-6884-6737;
Email: pdev@lps.umd.edu
Authors
Cyrille Armel Sayou Ngomsi – Department of Physics and Astronomy, Howard University, Washington, D.C. 20059, USA, Email: sayarmelo@gmail.com
Sai Krishna Narayanan – Department of Physics, University of Maryland, College Park, Maryland 20742, USA, Email: saik98@umd.edu
Author Contributions
PD conceived and directed the study. CASN and PD performed the DFT-based calculations. SKN calculated the defect formation energies with supercell-size corrections. PD wrote the final manuscript with inputs from CASN and SKN. All authors proof-read and reviewed the final revision of the manuscript.
Notes
The authors declare no competing financial interest.
@@ 3 Author Information
The research performed at the Howard University was supported by the National Science Foundation under NSF Grants Nos DMR-1738076 and OAC-2118099. This work used the Expanse and Bridges2 clusters at SDSC and PSC, respectively, through allocation PHY180014 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants No. 2138259, No. 2138286, No. 2138307, No. 2137603, and No. 2138296.